OpenStudy (anonymous):

Find the limit as h approaches 0, lim h/sin3h?

myininaya (myininaya):

Recall \[\lim_{u \rightarrow 0}\frac{u}{\sin(u)}=1\] By squeeze thm.

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{h}{\sin(3h)}=\lim_{h \rightarrow 0}\frac{a}{a} \frac{h}{\sin(3h)}\] \[\lim_{h \rightarrow 0}\frac{1}{a} \frac{ah}{\sin(3h)}=\frac{1}{a}\lim_{h \rightarrow 0}\frac{ah}{\sin(3h)}\] \[=\frac{1}{a}(1)=\frac{1}{a}\] what does a need to be in order for this to be true?

OpenStudy (anonymous):

a needs a 3

myininaya (myininaya):

Right on! :)

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{ah}{\sin(ah)}=1 , a \neq 0\]

myininaya (myininaya):

I hope this was helpful. Do you have any questions?

OpenStudy (anonymous):

yes that was very helpful, would you know how to graph f(x) (1-x2, x doesn't equal 1) (2, x=1)

myininaya (myininaya):

Ok so do you know how to just graph y=1-x^2 ?

OpenStudy (anonymous):

yes